What mass of NA2SO4. 7H2O contain exactly 6.022 *10^22 atoms of oxygen 

Grade:11

1 Answers

Arun
5 years ago
Dear Sachin
 
6.022* 10^22 * 1 molecues NA2SO4. 7H2O / 11 Oxygen atoms
 
 = 0.547 * 10^22 molecules NA2SO4. 7H2O 
 
Now
 
moles = 0.547 * 10^22 / 6.023 * 10^23 = 0.009 moles of NA2SO4. 7H2O 
 
hence mass  = 0.009 * (46+32+64+126) = 0.009 * 268 = 2.412 gm